### v3

# A note on the Björner–Kalai theorem

Abstract
In 1988, Björner and Kalai used combinatorial shadow functions to characterize the maximal Betti sequence for a given $f$-vector and the minimal $f$-vector for a given Betti sequence. Their description of the maximal Betti sequence was expressed through a set of inequalities. In this paper, we introduce an error function $\delta_{k}$ associated with the combinatorial shadow functions and use it to sharpen these inequalities into exact equalities. As a corollary, we obtain an equivalent form of Björner and Kalai’s characterization of all possible pairs $(f,\beta)$ that can occur as the $f$-vector and Betti sequence of a simplicial complex. Moreover, combining our results with a previous result of Björner in 2011, we derive a new number-theoretic inequality concerning the count of odd square-free integers with a specified number of prime factors.
Keywords: combinatorial shadow function, simplicial complex, $f$-vector, Betti number
2010 MSC: 05E45
1 Introduction
Let $k\geq 1$ be a fixed integer. According to number theory, for any integer $n\geq 1$, there exists a unique sequence of integers $a_{k}>a_{k-1}>\cdots>a_{i}\geq i\geq 1$ such that
$n=\binom{a_{k}}{k}+\binom{a_{k-1}}{k-1}+\cdots+\binom{a_{i}}{i}$. (1)
This unique expansion allows one to define the following combinatorial shadow functions:
$\partial_{k-1}(n)$ $=\binom{a_{k}}{k-1}+\binom{a_{k-1}}{k-2}+\cdots+\binom{a_{i}}{i-1}$,
$\partial^{k-1}(n)$ $=\binom{a_{k}-1}{k}+\binom{a_{k-1}-1}{k-1}+\cdots+\binom{a_{i}-1}{i}$.
For completeness, let $\partial_{k-1}(0)=\partial^{k-1}(0)=0$.
Let $\mathbb{N}_{0}^{(\infty)}$ denote the set of ultimately vanishing sequences of nonnegative integers. The set $\mathbb{N}_{0}^{(\infty)}$ and all its subsets are naturally equipped with the componentwise partial order, defined by
$(n_{0},n_{1},\ldots)\leq(n_{0}^{\prime},n_{1}^{\prime},\ldots)\quad\Longleftrightarrow\quad n_{i}\leq n_{i}^{\prime}\text{ for all }i\geq 0.$
For an ordered pair $(f=(f_{0},f_{1},\ldots),\beta=(\beta_{0},\beta_{1},\beta_{2},\ldots))$ of sequences from $\mathbb{N}_{0}^{(\infty)}$, we define
$\chi_{k-1}:=\chi_{k-1}(f,\beta)=\sum_{j\geq k}(-1)^{j-k}(f_{j}-\beta_{j}),\quad\text{for each }k\geq 0.$
Clearly, $f_{k}-\chi_{k-1}=\chi_{k}+\beta_{k}$.
Let $\Delta$ be a finite simplicial complex. The $f$-vector of $\Delta$ is the sequence $f=(f_{0},f_{1},\ldots)$, where
$f_{i}=\#\{F\in\Delta\mid\dim F=i\},\quad\text{for each }i\geq 0.$
Let $\boldsymbol{k}$ be a fixed field, and let $\widetilde{H}_{i}(\Delta;\boldsymbol{k})$ denote the $i$-th reduced simplicial homology group of $\Delta$ over $\boldsymbol{k}$. The reduced Betti sequence of $\Delta$ over $\boldsymbol{k}$ is defined as $\beta=(\beta_{0},\beta_{1},\ldots)$, where
$\beta_{i}=\dim_{\boldsymbol{k}}\widetilde{H}_{i}(\Delta;\boldsymbol{k}),\quad\text{for each }i\geq 0.$
Using the combinatorial shadow function $\partial_{k}$, Björner and Kalai [2] provided a complete characterization of all possible pairs $(f,\beta)$ that may occur as the $f$-vector and Betti sequence of a finite simplicial complex, extending the well-known Euler–Poincaré theorem (cf. [6, 7]):
$\chi_{-1}=\sum_{j\geq 0}(-1)^{j}(f_{j}-\beta_{j})=1.$
Theorem 1.1 (Björner and Kalai, [2, Theorem 1.1]).
Suppose that $f=(f_{0},f_{1},\ldots),\beta=(\beta_{0},\beta_{1},\ldots)\in\mathbb{N}_{0}^{(\infty)}$ are two given sequences and $\boldsymbol{k}$ is a field. Then the following conditions are equivalent:
  1. $(a)$
    $f$ is the $f$-vector and $\beta$ the Betti sequence over $\boldsymbol{k}$ of some simplicial complex,
  2. $(b)$
    $\chi_{-1}=1$, and $\partial_{k}(\chi_{k}+\beta_{k})\leq\chi_{k-1}$ for all $k\geq 1$.
An ordered pair $(f,\beta)$ of sequences from $\mathbb{N}_{0}^{(\infty)}$ is called compatible, denoted by $f\sim\beta$, if there exists a simplicial complex with $f$-vector $f$ and Betti sequence $\beta$. By Theorem 1.1, this relation is purely combinatorial and independent of the field characteristic.
In their seminal work [2], Björner and Kalai further established the following foundational result concerning compatible pairs $(f,\beta)$.
Theorem 1.2 (Björner and Kalai, [2, Theorem 1.2, Theorem 5.3]).
  1. $(a)$
    Suppose that $f$ is the $f$-vector of some simplicial complex. Then the set $B_{f}=\{\beta\in\mathbb{N}_{0}^{(\infty)}:f\sim\beta\}$ has a unique maximal element. Define $\psi(f)=\max B_{f}$.
  2. $(b)$
    Suppose that $\beta\in\mathbb{N}_{0}^{(\infty)}$. The set $F_{\beta}=\{f\in\mathbb{N}_{0}^{(\infty)}:f\sim\beta\}$ has a unique minimal element. Define $\phi(\beta)=\min F_{\beta}$.
  3. $(c)$
    Suppose that $f$ is the $f$-vector of some simplicial complex, and $\beta\in\mathbb{N}_{0}^{(\infty)}$. Then $\beta=\psi(f)$ if and only if $\chi_{-1}=1$, and $\partial_{k}(\chi_{k}+\beta_{k})\leq\chi_{k-1}\leq\partial_{k}(\chi_{k}+\beta_{k}+1)$ for all $k\geq 1$.
  4. $(d)$
    Suppose that $(f,\beta)$ is a compatible pair. Then $f=\phi(\beta)$ if and only if $\partial_{k}(\chi_{k}+\beta_{k})=\chi_{k-1}$ for all $k\geq 1$.
  5. $(e)$
    $\psi(\phi(\beta))=\beta$ for all $\beta\in\mathbb{N}_{0}^{(\infty)}$.
  6. $(f)$
    $\phi(\psi(f))\leq f$ for all $f$-vectors $f$.
We note that only the maximal part contains inequalities, which inspired us to introduce a new function to measure the discrepancy in these inequalities. The error function $\delta_{k-1}$ is defined by
$\delta_{k-1}(n)=\ell,$
where $\ell$ is the number of indices $t$ in the unique expansion (1) for which $a_{t}=t$. For completeness, we define $\delta_{k-1}(0)=0$.
In this paper, we prove the following result, which shows how the error function $\delta_{k}$ affects inequalities related to the combinatorial shadow functions $\partial_{k}$ and $\partial^{k}$.
Theorem 1.3.
Let $n$, $m$ be nonnegative integers and let $\epsilon_{k}$ be an integer with $0\leq\epsilon_{k}\leq\delta_{k}(n+m)$. Then the following statements hold:
  1. $(a)$
    $\partial_{k}(n)\leq m-\epsilon_{k}$ if and only if $n\leq\partial^{k}(n+m)$;
  2. $(b)$
    $n=\partial^{k}(n+m)$ if and only if $\partial_{k}(n)+\delta_{k}(n+m)=m$;
  3. $(c)$
    $\partial_{k}(n)=m$ if and only if $n=\partial^{k}(n+m)$ and $\delta_{k}(n+m)=0$.
As an application of Theorem 1.3, we sharpen the inequalities in parts $(c)$ and $(f)$ of Theorem 1.2 and provide an equivalent reformulation of part $(d)$. For a sequence $f=(f_{0},f_{1},\ldots)\in\mathbb{N}_{0}^{(\infty)}$, we define
$\delta(f)=(0,\delta_{1}(f_{1}),\delta_{2}(f_{2}),\ldots)\quad\text{and}\quad\delta_{+}(f)=(\delta_{1}(f_{1}),\delta_{2}(f_{2}),\ldots).$
Theorem 1.4.
  1. $(a)$
    Suppose that $f$ is the $f$-vector of some simplicial complex. Then the set $B_{f}=\{\beta\in\mathbb{N}_{0}^{(\infty)}:f\sim\beta\}$ has a unique maximal element. Define $\psi(f)=\max B_{f}$.
  2. $(b)$
    Suppose that $\beta\in\mathbb{N}_{0}^{(\infty)}$. The set $F_{\beta}=\{f\in\mathbb{N}_{0}^{(\infty)}:f\sim\beta\}$ has a unique minimal element. Define $\phi(\beta)=\min F_{\beta}$.
  3. $(c^{*})$
    Suppose that $f$ is the $f$-vector of some simplicial complex, and $\beta\in\mathbb{N}_{0}^{(\infty)}$. Then $\beta=\psi(f)$ if and only if $\chi_{-1}=1$, and $\partial_{k}(\chi_{k}+\beta_{k})+\delta_{k}(f_{k})=\chi_{k-1}$ for all $k\geq 1$, which is the case if and only if $\chi_{-1}=1$, and $f_{k}-\chi_{k-1}=\partial^{k}(f_{k})$ for all $k\geq 1$.
  4. $(d^{*})$
    Suppose that $(f,\beta)$ is a compatible pair. Then $f=\phi(\beta)$ if and only if $\partial_{k}(\chi_{k}+\beta_{k})=\chi_{k-1}$ for all $k\geq 1$, which is the case if and only if $f_{k}-\chi_{k-1}=\partial^{k}(f_{k})$ for all $k\geq 1$, and $\delta(f)=0$.
  5. $(e)$
    $\psi(\phi(\beta))=\beta$ for all $\beta\in\mathbb{N}_{0}^{(\infty)}$.
  6. $(f^{*})$
    $\phi(\psi(f)+\delta_{+}(f))=f-\delta(f)$ for all $f$-vectors $f$. Therefore, $\phi(\psi(f))\leq f-\delta(f)-\delta_{+}(f)$, with equality holding if and only if $\delta(f)=0$.
By applying Theorem 1.3 to Theorem 1.1, we also obtain the following result.
Corollary 1.5.
Suppose that $f=(f_{0},f_{1},\ldots),\beta=(\beta_{0},\beta_{1},\ldots)\in\mathbb{N}_{0}^{(\infty)}$ are two given sequences and $\boldsymbol{k}$ is a field. Let $\epsilon\in\mathbb{N}_{0}^{(\infty)}$ be any sequence satisfying $0\leq\epsilon\leq\delta(f)$. Then the following conditions are equivalent:
  1. $(a)$
    $f$ is the $f$-vector and $\beta$ the Betti sequence over $\boldsymbol{k}$ of some simplicial complex,
  2. $(b^{*})$
    $\chi_{-1}=1$, and $\partial_{k}(\chi_{k}+\beta_{k})\leq\chi_{k-1}-\epsilon_{k}$ for all $k\geq 1$,
  3. $(b^{+})$
    $\chi_{-1}=1$, and $0\leq f_{k}-\chi_{k-1}\leq\partial^{k}(f_{k})$ for all $k\geq 1$.
Remark 1.6.
It should be noted that condition $(b^{*})$ in Corollary 1.5 provides a refinement of condition $(b)$ from Theorem 1.1, while condition $(b^{+})$ establishes a new characterization of compatible pairs $(f,\beta)$. Furthermore, for any simplicial complex, the equality
$f_{k}-\chi_{k-1}=\dim Z_{k}$
holds, where $Z_{k}$ represents the space of $k$-cycles (see [2, Remark 4.4]). Therefore, through homological interpretation of the inequality $f_{k}-\chi_{k-1}\leq\partial^{k}(f_{k})$ in Corollary 1.5 $(b^{+})$, we obtain
$\dim Z_{k}\leq\partial^{k}(f_{k}).$
Let $m$ be a square-free positive integer, and let $P(m)$ be the set of prime factors of $m$. The family
$\Delta_{n}:=\{P(m):m\text{ is square-free and }m\leq n\}$
is a simplicial complex, called the Björner complex. It was shown by Björner [1] that the Prime Number Theorem and the Riemann Hypothesis are equivalent to certain asymptotic estimates of the reduced Euler characteristics of these complexes as $n\to\infty$.
Let $\sigma_{k}(n)$ (resp. $\sigma_{k}^{\mathrm{odd}}(n)$) denote the number of square-free integers (resp. odd square-free integers) $\leq n$ with exactly $k$ prime factors. In [1], Björner applied Theorem 1.1 to the Björner complex and proved that
$\partial_{k}(\sigma_{k+1}^{\mathrm{odd}}(n))\leq\sigma_{k}^{\mathrm{odd}}(n/2)\quad\text{for all }k\geq 1.$
Combining this result with Theorem 1.3, we immediately derive the following number-theoretic inequality.
Corollary 1.7.
For all $k\geq 1$,
$\sigma_{k+1}^{\mathrm{odd}}(n)\leq\partial^{k}(\sigma_{k+1}(n)).$
2 Proofs of Theorems 1.3 and 1.4
In order to prove Theorem 1.3, we need a series of number-theoretic lemmas.
Lemma 2.1 ([2], [3, Lemma 3.1]).
Let $n$ and $k$ be positive integers. Then there is a unique expansion
$n=\binom{a_{k}}{k}+\binom{a_{k-1}}{k-1}+\cdots+\binom{a_{i}}{i}$,
such that $a_{k}>a_{k-1}>\cdots>a_{i}\geq i\geq 1$. This expansion is called the (k1)(k-1)-dimensional representation of $n$.
Lemma 2.2.
Let $n$, $m$, and $k$ be positive integers. Suppose that
$n=\binom{a_{k+1}}{k+1}+\binom{a_{k}}{k}+\cdots+\binom{a_{i}}{i}$
and
$m=\binom{b_{k+1}}{k+1}+\binom{b_{k}}{k}+\cdots+\binom{b_{j}}{j}$
are the $k$-dimensional representations of $n$ and $m$, respectively. Then the following statements hold:
  1. $(a)$
    $n=m$ if and only if $i=j$ and $a_{t}=b_{t}$ for all $t\in\{i,\ldots,k+1\}$;
  2. $(b)$
    $n<m$ if and only if either $i>j$ and $a_{t}=b_{t}$ for all $t\in\{i,\ldots,k+1\}$, or the maximal index $t\in\{\max\{i,j\},\ldots,k+1\}$ with $a_{t}\neq b_{t}$ satisfies $a_{t}<b_{t}$.
Proof.
Note that $(a)$ follows from Lemma 2.1 immediately. Thus it suffices to consider $(b)$. Using Pascal’s identity $\binom{p}{q}=\binom{p-1}{q}+\binom{p-1}{q-1}$ recursively, we obtain
$\binom{c+1}{d}=\binom{c}{d}+\binom{c-1}{d-1}+\cdots+\binom{c-d}{0}$
for any integers $0\leq d\leq c+1$. This implies that
$\binom{c_{p}}{p}>\binom{d_{p}}{p}+\binom{d_{p-1}}{p-1}+\cdots+\binom{d_{l}}{l}$ (2)
whenever $c_{p}>d_{p}$ and $d_{p}>d_{p-1}>\cdots>d_{l}\geq l\geq 1$. Let $s=\max\{i,j\}$. If $a_{t}=b_{t}$ for all $t\in\{s,\ldots,k+1\}$, then $n<m$ when $i>j$, $n>m$ when $i<j$, and $n=m$ when $i=j$. Now suppose that there exists some index $t\in\{s,\ldots,k+1\}$ such that $a_{t}\neq b_{t}$. Let $t_{0}=\max\{t\in\{s,\ldots,k+1\}:a_{t}\neq b_{t}\}$. Then it follows from (2) that $n<m$ when $a_{t_{0}}<b_{t_{0}}$, and $n>m$ when $a_{t_{0}}>b_{t_{0}}$. This proves $(b)$. ∎
For any integer $n,k\geq 1$, using the unique expansion (1), we define:
$d^{k-1}(n)=\binom{a_{k}+1}{k}+\binom{a_{k-1}+1}{k-1}+\cdots+\binom{a_{i}+1}{i}$.
For completeness, define $d^{k-1}(0)=0$.
Lemma 2.3.
Let $n$ be a nonnegative integer. Then $n+\partial_{k}(n)=d^{k}(n)$.
Proof.
The case $n=0$ is trivial, so we may assume $n\geq 1$. By Lemma 2.1, we suppose that the $k$-dimensional representation of $n$ is given by
$n=\binom{a_{k+1}}{k+1}+\binom{a_{k}}{k}+\cdots+\binom{a_{i}}{i}$,
where $a_{k+1}>a_{k}>\cdots>a_{i}\geq i\geq 1$. Thus we have
$n+\partial_{k}(n)$ $=\binom{a_{k+1}}{k+1}+\binom{a_{k}}{k}+\cdots+\binom{a_{i}}{i}+\binom{a_{k+1}}{k}+\binom{a_{k}}{k-1}+\cdots+\binom{a_{i}}{i-1}$
$=\binom{a_{k+1}+1}{k+1}+\binom{a_{k}+1}{k}+\cdots+\binom{a_{i}+1}{i}$
$=d^{k}(n),$
as desired. ∎
Lemma 2.4.
Let $n$ be a nonnegative integer. Then $\delta_{k}(n)=n-\partial^{k}(n)-\partial_{k}(\partial^{k}(n))$.
Proof.
The case $n=0$ is trivial, so we may assume $n\geq 1$. By Lemma 2.1, we suppose that the $k$-dimensional representation of $n$ is given by
$n=\binom{a_{k+1}}{k+1}+\binom{a_{k}}{k}+\cdots+\binom{a_{i}}{i}$,
where $a_{k+1}>a_{k}>\cdots>a_{i}\geq i\geq 1$. If $a_{t}=t$ for all $t\in\{i,\ldots,k+1\}$, then $\delta_{k}(n)=n=k+2-i$, $\partial^{k}(n)=\partial_{k}(\partial^{k}(n))=0$, and we are done. Now suppose that there exists some $t\in\{i,\ldots,k+1\}$ such that $a_{t}\neq t$. Let $t_{0}=\min\{t\in\{i,\ldots,k+1\}:a_{t}\neq t\}$. Then $a_{t}=t$ when $t<t_{0}$, and $a_{t}>t$ when $t\geq t_{0}$. According to the definition of the operator $\delta_{k}$, we obtain
$\delta_{k}(n)=t_{0}-i=\binom{a_{t_{0}-1}}{t_{0}-1}+\cdots+\binom{a_{i}}{i}$.
Therefore, by Lemma 2.3,
$\partial^{k}(n)+\partial_{k}(\partial^{k}(n))$ $=d^{k}(\partial^{k}(n))$
$=d^{k}\left(\binom{a_{k+1}-1}{k+1}+\binom{a_{k}-1}{k}+\cdots+\binom{a_{i_{0}}-1}{i_{0}}\right)$
$=\binom{a_{k+1}}{k+1}+\binom{a_{k}}{k}+\cdots+\binom{a_{i_{0}}}{i_{0}}$
$=n-\delta(n),$
as desired. ∎
Lemma 2.5.
Let $n$ and $m$ be nonnegative integers.
  1. $(a)$
    If $d^{k}(n)\leq m$, then $n\leq\partial^{k}(m)$. In particular, if $d^{k}(n)=m$, then $n=\partial^{k}(m)$.
  2. $(b)$
    If $n\leq m$, then $\partial_{k}(n)\leq\partial_{k}(m)$.
  3. $(c)$
    If $d^{k}(n)=d^{k}(m)$, then $n=m$.
Proof.
If $n=0$ or $m=0$, the result is trivial. Thus we may assume that $n,m\geq 1$. By Lemma 2.1, we suppose that the $k$-dimensional representations of $n$ and $m$ are given by
$n=\binom{a_{k+1}}{k+1}+\binom{a_{k}}{k}+\cdots+\binom{a_{i}}{i}$
and
$m=\binom{b_{k+1}}{k+1}+\binom{b_{k}}{k}+\cdots+\binom{b_{j}}{j}$,
where $a_{k+1}>a_{k}>\cdots>a_{i}\geq i\geq 1$ and $b_{k+1}>b_{k}>\cdots>b_{j}\geq j\geq 1$, respectively. Then
$d^{k}(n)=\binom{a_{k+1}+1}{k+1}+\binom{a_{k}+1}{k}+\cdots+\binom{a_{i}+1}{i}$
and
$d^{k}(m)=\binom{b_{k+1}+1}{k+1}+\binom{b_{k}+1}{k}+\cdots+\binom{b_{j}+1}{j}$
are the $k$-dimensional representations of $d^{k}(n)$ and $d^{k}(m)$, respectively.
First consider $(a)$. If $d^{k}(n)=m$, by Lemma 2.2, we obtain $i=j$ and $a_{t}+1=b_{t}$ for all $t\in\{i,\ldots,k+1\}$. Therefore, $n=\partial^{k}(m)$. If $d^{k}(n)<m$, again by Lemma 2.2, either $i>j$ and $a_{t}+1=b_{t}$ for all $t\in\{i,\ldots,k+1\}$, or the maximal index $t\in\{\max\{i,j\},\ldots,k+1\}$ with $a_{t}+1\neq b_{t}$ satisfies $a_{t}+1<b_{t}$. For the former case, we have
$n$ $=\binom{a_{k+1}}{k+1}+\binom{a_{k}}{k}+\cdots+\binom{a_{i}}{i}$
$=\binom{b_{k+1}-1}{k+1}+\binom{b_{k}-1}{k}+\cdots+\binom{b_{i}-1}{i}$
$\leq\binom{b_{k+1}-1}{k+1}+\binom{b_{k}-1}{k}+\cdots+\binom{b_{i}-1}{i}+\cdots+\binom{b_{j}-1}{j}$
$=\partial^{k}(m),$
as desired. For the latter case, let $t_{0}=\max\{t\in\{\max\{i,j\},\ldots,k+1\}:a_{t}+1\neq b_{t}\}$. Then $a_{t_{0}}+1<b_{t_{0}}$, and we have
$n$ $=\binom{a_{k+1}}{k+1}+\binom{a_{k}}{k}+\cdots+\binom{a_{i}}{i}$
$<\binom{a_{k+1}}{k+1}+\binom{a_{k}}{k}+\cdots+\binom{a_{t_{0}+1}}{t_{0}+1}+\binom{b_{t_{0}}-1}{t_{0}}$
$=\binom{b_{k+1}-1}{k+1}+\binom{b_{k}-1}{k}+\cdots+\binom{b_{t_{0}+1}-1}{t_{0}+1}+\binom{b_{t_{0}}-1}{t_{0}}$
$\leq\binom{b_{k+1}-1}{k+1}+\binom{b_{k}-1}{k}+\cdots+\binom{b_{j}-1}{j}$
$=\partial^{k}(m).$
This proves $(a)$.
Now consider $(b)$. If $n=m$, then $\partial_{k}(n)=\partial_{k}(m)$ by definition. Thus we can assume that $n<m$. Again by Lemma 2.2, we have that either $i>j$ and $a_{t}=b_{t}$ for all $t\in\{i,\ldots,k+1\}$, or the maximal index $t\in\{\max\{i,j\},\ldots,k+1\}$ with $a_{t}\neq b_{t}$ satisfies $a_{t}<b_{t}$. For the former case, we have
$\partial_{k}(n)$ $=\binom{a_{k+1}}{k}+\binom{a_{k}}{k-1}+\cdots+\binom{a_{i}}{i-1}$
$=\binom{b_{k+1}}{k}+\binom{b_{k}}{k-1}+\cdots+\binom{b_{i}}{i-1}$
$\leq\binom{b_{k+1}}{k}+\binom{b_{k}}{k-1}+\cdots+\binom{b_{i}}{i-1}+\cdots+\binom{b_{j}}{j-1}$
$=\partial_{k}(m).$
For the latter case, let $t_{0}=\max\{t\in\{\max\{i,j\},\ldots,k+1\}:a_{t}\neq b_{t}\}$. Then $a_{t_{0}}<b_{t_{0}}$, and we obtain
$\partial_{k}(n)-\partial_{k}(m)$ $=\binom{a_{t_{0}}}{t_{0}-1}+\cdots+\binom{a_{i}}{i-1}-\binom{b_{t_{0}}}{t_{0}-1}-\cdots-\binom{b_{j}}{j-1}$
$\leq\binom{a_{t_{0}}}{t_{0}-1}+\cdots+\binom{a_{i}}{i-1}-\binom{b_{t_{0}}}{t_{0}-1}$
$\leq\binom{a_{t_{0}}}{t_{0}-1}+\cdots+\binom{a_{i}}{i-1}-\binom{a_{t_{0}}+1}{t_{0}-1}$
$\leq\binom{a_{t_{0}}}{t_{0}-1}+\binom{a_{t_{0}}-1}{t_{0}-2}+\cdots+\binom{a_{t_{0}}-t_{0}+1}{0}-\binom{a_{t_{0}}+1}{t_{0}-1}$
$=\binom{a_{t_{0}}+1}{t_{0}-1}-\binom{a_{t_{0}}+1}{t_{0}-1}$
$=0,$
where the penultimate equality follows from Pascal’s identity. Thus $(b)$ follows.
Next consider $(c)$. If $d^{k}(n)=d^{k}(m)$, by Lemma 2.2, we assert that $i=j$ and $a_{t}+1=b_{t}+1$ for all $t\in\{i,\ldots,k+1\}$. Therefore, $n=m$, and $(c)$ follows. ∎
We are now ready to give the proof of Theorem 1.3.
Proof of Theorem 1.3.
If $\partial_{k}(n)\leq m-\epsilon_{k}$, by Lemma 2.3,
$d^{k}(n)=n+\partial_{k}(n)\leq n+m-\epsilon_{k}.$
Since $0\leq\epsilon_{k}\leq\delta_{k}(f_{k})$, by Lemma 2.5 $(a)$, we deduce that
$n\leq\partial^{k}(n+m-\epsilon_{k})=\partial^{k}(n+m).$
Conversely, if $n\leq\partial^{k}(n+m)$, by Lemma 2.5 $(b)$, we have
$\partial_{k}(n)\leq\partial_{k}(\partial^{k}(n+m)),$
and hence
$n+\partial_{k}(n)\leq\partial^{k}(n+m)+\partial_{k}(\partial^{k}(n+m)).$
Combining this with Lemma 2.4 yields that
$\partial_{k}(n)$ $\leq m-(n+m-\partial^{k}(n+m)-\partial_{k}(\partial^{k}(n+m)))=m-\delta_{k}(n+m)\leq m-\epsilon_{k}.$
This proves $(a)$.
If $n=\partial^{k}(n+m)$, then $\partial_{k}(n)=\partial_{k}(\partial^{k}(n+m))$. This implies
$n+\partial_{k}(n)=\partial^{k}(n+m)+\partial_{k}(\partial^{k}(n+m)),$ (3)
or equivalently,
$\partial_{k}(n)+\delta_{k}(n+m)=m$
by Lemma 2.4. Conversely, if $\partial_{k}(n)+\delta_{k}(n+m)=m$, then we recover (3). By Lemma 2.3, it follows that
$d^{k}(n)=d^{k}(\partial^{k}(n+m)).$
Combining this with Lemma 2.5 $(c)$, we obtain
$n=\partial^{k}(n+m).$
Thus $(b)$ follows.
Suppose that $n=\partial^{k}(n+m)$. According to $(b)$, we have $\partial_{k}(n)+\delta_{k}(n+m)=m$. If $\delta_{k}(n+m)=0$ then $\partial_{k}(n)=m$, and if $\delta_{k}(n+m)\neq 0$, then $\partial_{k}(n)\neq m$. On the other hand, if $\partial_{k}(n)=m$, by Lemma 2.3,
$d^{k}(n)=\partial_{k}(n)+n=m+n.$
Combining this with Lemma 2.5 $(a)$, we obtain $n=\partial^{k}(n+m)$, and so $(c)$ follows.
This completes the proof. ∎
In order to prove Theorem 1.4, we also need the Kruskal-Katona theorem [4, 5], which gives a complete characterization of $f$-vectors of simplicial complexes.
Lemma 2.6 (Kruskal [5], Katona [4]).
A sequence $f=(f_{0},f_{1},\ldots)\in\mathbb{N}_{0}^{(\infty)}$ is the $f$-vector of some simplicial complex if and only if
$\partial_{k}(f_{k})\leq f_{k-1},\quad\text{for every}~k\geq 1.$
In what follows, we present the proof of Theorem 1.4.
Proof of Theorem 1.4.
First observe that $(d^{*})$ follows from Theorem 1.2 and Theorem 1.3 immediately. Now suppose that $f$ is the $f$-vector of some simplicial complex. Let $\beta\in B_{f}$. Since $f\sim\beta$, by Corollary 1.5, we have $\chi_{-1}=1$, and $f_{k}-\chi_{k-1}\leq\partial^{k}(f_{k})$ for all $k\geq 1$. Also note that $f_{0}-\chi_{-1}=f_{0}-1=\partial^{0}(f_{0})$. Hence, for each $k\geq 0$,
$\beta_{k}$ $=f_{k}-\chi_{k}-\chi_{k-1}$ (4)
$=(f_{k}-\chi_{k-1})+(f_{k+1}-\chi_{k})-f_{k+1}$
$\leq\partial^{k}(f_{k})+\partial^{k+1}(f_{k+1})-f_{k+1}.$
Furthermore, according to [2, Theorem 5.3], the upper bound in (4) can be achieved uniformly, that is,
$\psi(f)=\max B_{f}=(\psi_{0},\psi_{1},\ldots)~\text{with}~\psi_{i}=\partial^{i}(f_{i})+\partial^{i+1}(f_{i+1})-f_{i+1}~\text{for}~i\geq 0.$
In order to prove $(c^{*})$, by Theorem 1.2 and Theorem 1.3, it suffices to prove that
$\beta=\psi(f)\Longleftrightarrow\chi_{-1}=1~\text{and}~f_{k}-\chi_{k-1}=\partial^{k}(f_{k})~\text{for all}~k\geq 1.$
If $\beta=\psi(f)$, then $f\sim\beta$, and hence $\chi_{-1}=1$. Moreover, from (4) we obtain $f_{k}-\chi_{k-1}=\partial^{k}(f_{k})$ for all $k\geq 1$. Conversely, if $\chi_{-1}=1$ and $f_{k}-\chi_{k-1}=\partial^{k}(f_{k})$ for all $k\geq 1$, then
$\beta_{k}=f_{k}-\chi_{k}-\chi_{k-1}=f_{k}-\chi_{k-1}+f_{k+1}-\chi_{k}-f_{k+1}=\partial^{k}(f_{k})+\partial^{k+1}(f_{k+1})-f_{k+1}$
for all $k\geq 0$. Therefore, $\beta=\psi(f)$. This proves $(c^{*})$.
Next we consider $(f^{*})$. According to $(c^{*})$ and $(d^{*})$, if $f$ is an $f$-vector with $\delta(f)=0$, then $\phi(\psi(f))=f$. Now suppose that $f$ is an arbitrary $f$-vector. By Lemma 2.6, it is easy to see that $f-\delta(f)$ is also an $f$-vector. Moreover, by definition, we have $\delta(f-\delta(f))=0$. Thus $\phi(\psi(f-\delta(f)))=f-\delta(f)$ by the above arguments. Note that, for any $k\geq 0$,
$\psi_{k}(f_{k}-\delta_{k}(f_{k}))$ $=\partial^{k}(f_{k}-\delta_{k}(f_{k}))+\partial^{k+1}(f_{k+1}-\delta_{k+1}(f_{k+1}))-f_{k+1}+\delta_{k+1}(f_{k+1})$
$=\partial^{k}(f_{k})+\partial^{k+1}(f_{k+1})-f_{k+1}+\delta_{k+1}(f_{k+1})$
$=\psi_{k}(f_{k})+\delta_{k+1}(f_{k+1}).$
This implies $\psi(f-\delta(f))=\psi(f)+\delta_{+}(f)$, and consequently,
$\phi(\psi(f)+\delta_{+}(f))=\phi(\psi(f-\delta(f)))=f-\delta(f).$
By [2, Theorem 5.2], the mapping $\phi$ is injective and order-preserving. Therefore, we conclude that
$\phi(\psi(f))+\delta_{+}(f)\leq\phi(\psi(f)+\delta_{+}(f))=f-\delta(f),$
where equalities hold everywhere if and only if $\delta(f)=0$.
This completes the proof. ∎
Declaration of interest statement
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
X. Huang was supported by the National Natural Science Foundation of China (No. 12471324) and the Natural Science Foundation of Shanghai (No. 24ZR1415500).
Data availability
No data was used for the research described in the article.